### v1

# Generalized Landau Yang Theorem

Abstract
Landau Yang theorem is well known for the past several decades. It prohibits decay of a massive spin 1 particle to two photons. This emerges simply from the representation theory of the Poincare group and Bose Statistics. It does not require any action or Lagrangian. We generalize this theorem to theories with supersymmetry (SUSY) which disallows even decay to two photinos (Majorana fermions). We will prove that if the photon has a mass, howsoever small, this theorem can be evaded. We also show that the supersymmetric selection rule above can also be evaded through the Stueckelberg mass term. Further interesting implications are also pointed out.
Generalized Landau Yang Theorem
T. R. Govindarajan111trg@imsc.res.in, govindarajan.thupil@krea.edu.in
The Institute of Mathematical Sciences, Chennai 600113, Tamil Nadu, India
Krea University, Sricity, 517646, Andhra Pradesh, India
Rakesh Tibrewala222rtibs@lnmiit.ac.in
The LNMIIT, Jaipur 302031, Rajasthan, India
October 07, 2024
1 Introduction:
Landau[1] and Yang [2] independently proved a theorem which disallows a massive spin 1 particle from decaying to two massless spin-1 bosons without further internal symmetries. The elegance of the theorem stems from the fact that it does not depend on any interactions and is very generic. This is easy to understand from the representation theory of the Poincare group and Bose-Einstein statistics. For example, the $\rho$ meson which has spin 1 does not decay to two photons whereas the pion (which has spin 0) does. It prohibits the eletroweak $Z$ particle to decay to two photons (the bound on the branching ratio $\Gamma_{Z\rightarrow\gamma\gamma}/\Gamma_{\mathrm{Tot}}<5.2\times 10^{-5}$ [3]). Recently this theorem has been applied to understand that the Higgs Boson cannot be a spin 1 particle due to its decay to two photons [4, 5].
1.1 Simple proof
Massive spin 1 particle can be taken to be at rest with the four momentum $P^{\mu}~{}=~{}(m,0,0,0)$. Its spin can be given as $\vec{S}$. The massless photons will have momenta $\vec{k}$ and $-\vec{k}$ in order to conserve the momentum. The polarization of these two photons will be in the plane perpendicular to $\vec{k}$. We can call them as $\vec{s}_{1},\vec{s}_{2}$. They will obey
$\vec{k}\cdot\vec{s}_{1}~{}=~{}0~{}=~{}\vec{k}\cdot\vec{s}_{2}$ (1)
Now let us write down all possible terms linear in $\vec{S},\vec{s}_{1},\vec{s}_{2}$ that can appear in the amplitude for the decay of the massive spin 1 particle to two photons. There are only three generic terms that can be written as candidates for the amplitudes. These are:
  1. 1.
    $A_{1}~{}=~{}a_{1}(k^{2})\vec{S}\cdot\vec{k}~{}\vec{s}_{1}\cdot\vec{s}_{2}$
  2. 2.
    $A_{2}~{}=~{}a_{2}(k^{2})(\vec{s}_{1}\times\vec{s}_{2})~{}\cdot~{}\vec{S}$
  3. 3.
    $A_{3}~{}=~{}a_{3}(k^{2})(\vec{s}_{1}\times\vec{s}_{2}~{}\cdot\vec{k})~{}(\vec{S}\cdot\vec{k})$
The general form of the amplitude is
${\cal{A}}(\vec{k},\vec{s}_{1},\vec{s}_{2},\vec{S})~{}=~{}\sum_{1}^{3}a_{i}(k^{2})~{}A_{i}$ (2)
Here $a_{i}(k^{2})$ are arbitrary invariant functions. Now the Bose symmetry of photons implies that the amplitude should be the same if we transform $\vec{k}\rightarrow-\vec{k}$ and interchange $\vec{s}_{1}~{}\leftrightarrow~{}\vec{s}_{2}$. That is we should have
${\cal{A}}(\vec{k},\vec{s}_{1},\vec{s}_{2},\vec{S})~{}=~{}{\cal{A}}(-\vec{k},\vec{s}_{2},\vec{s}_{1},\vec{S})$ (3)
But ${\cal{A}}(-\vec{k},\vec{s}_{2},\vec{s}_{1},\vec{S})~{}=~{}-{\cal{A}}(\vec{k},\vec{s}_{1},\vec{s}_{2},\vec{S})$ implying
${\cal{A}}(\vec{k},\vec{s}_{1},\vec{s}_{2},\vec{S})~{}=~{}0$ (4)
The vanishing of the amplitude is generic and does not depend on the details of the action or the nature of discrete transformation property of the initial vector particle. We can see that the above theorem stems only from the representation theory of the Poincare group and Bose symmetry.
1.2 Representation theory of Poincare symmetry
We recollect [6] that massless irreducible representations of the Poincare group are given by the helicity. On the other hand, massive representations are specified by the mass parameter as well as the spin value coming from the Pauli-Lubanski vector. We also need to represent the massless photons covariantly using a four vector $A_{\mu}$ to discuss further interactions through gauge theory. That brings in the helicity ±$\pm~{}1$. These follow from the induced representations of the little group of four momentum vector $p^{\mu}~{}=~{}(\omega,0,0,\omega)$ of a massless particle which is the Euclidean group in two dimensions. The little group for a massive particle on the other hand is $SO(3)$ which comes from the fact that the massive particle can be brought to rest in a frame and its four momentum in that frame is $(m,0,0,0)$. In addition it is well known that the helcity ±$\pm~{}1$ representations from $A_{\mu}$ are reducible but not decomposable[7]. We need to look at the content of the product of two such helicity representations through Clebsch Gordon coefficients. Also Bose symmetry should be imposed. It is clear that the decay of a massive spin 1 particle to two photons to be possible demands that the symmetric product of two massless representations of the Poincare group contains a massive spin one representation. It was shown by Balachandran and Jo[8] that it does not contain such a representation. The direct product of two massless representations with four momenta $(p,0,0,\pm p)$ and helicities $\lambda_{1},\lambda_{2}$ can be given by:
[…]
where $d\xi(R)$ is the Haar measure on $SU(2)$ and $D_{\mu~{}\lambda_{1}-\lambda_{2}}^{j}$ are the spin-j representation matrices. For the case of two identical particles we have
$|\lambda_{1}~{}\lambda_{2}~{}\hat{p}~{}j~{}\mu{\rangle}~{}=~{}\frac{1}{2}\left(|\lambda_{1}~{}\lambda_{2}~{}\hat{p}~{}j~{}\mu{\rangle}~{}\pm~{}(-1)^{j+\lambda_{1}+\lambda_{2}}~{}|\lambda_{2}~{}\lambda_{1}~{}\hat{p}~{}j~{}\mu{\rangle}~{}\right)$ (6)
For photons obeying Bose statistics we need the symmetric product. For the symmetric product with two identical particles of helicities $\lambda_{i}=\pm 1,i=1,2$ we have for $\lambda_{1}=\pm\lambda_{2}$:
$|\lambda_{1}~{}\lambda_{2}~{}\hat{p}~{}j~{}\mu{\rangle}~{}_{S}~{}=~{}\frac{1}{2}(1~{}+~{}(-1)^{j+\lambda_{1}+\lambda_{2}})~{}|\lambda_{1}~{}\lambda_{2}~{}\hat{p}~{}j~{}\mu{\rangle}~{}$ (7)
With $j=1$ we find the right hand side vanishes. This implies two photon states cannot have $j=1$ component thereby forbidding, for example, $Z\rightarrow 2\gamma$ transition and many others. On the other hand if $j=0$ as we expect for the Higgs, the decay to two photons is not prohibited by the theorem and provides a strong evidence for the scalar nature of Higgs [4, 5].
2 Generalization to SUSY
We will now study the extension of the above theorem in the Supersymmetric context. The obvious starting point is the representation theory of the Super Poincare algebra.
Super Poincare algebra also has two Casimir operators. The first is the usual $P_{\mu}~{}P^{\mu}$ corresponding to the mass of the particles. But the spin from the Pauli Lubanski vector does not provide the other invariant. There is a generalization of this, known as the superspin operator which commutes with all the generators of the super Poincare algebra [9]. The mass $m$ and the superspin operator $Y$ provide the massive irreducible representations. When Y = 0 we get a supersymmetric multiplet consisting of a scalar and a spin $1/2$ particle of the same mass.
The vector multiplet $V$ (obtained for $Y=1/2$) consists of a massive vector particle (Z), two chiral spinors (zinos) and a scalar ($\phi$). On the other hand, for the massless representation of photon $\gamma$ we have only the Majorana particle photino $\tilde{\gamma}$ as the fermionic partner.
Given the above particle content we can easily see the extension of the Landau Yang theorem to forbid decay of $Z\rightarrow\gamma+\gamma$. Not only that, in the effective field theory we will not have any interaction term representing the three superfields whose content includes $Z,\gamma,\gamma$ as external legs.
Hence combined with supersymmetry we will also have in addition zero amplitude for Z decaying to two photinos. Here the Fermi symmetry of photinos along with SUSY forbids such a process. Interestingly such a decay is prohibited by simply Poincare group representation theory itself – a massive spin 1 particle cannot decay to two (identical) massless Majorana particles, Fig. 1(a). This is seen in Balachandran and Jo paper through the decomposition of the antisymmetric product of two spin $1/2$ Majorana representations. To see this consider Eq. (7) with $j=1,\lambda_{i}=1/2$ corresponding to decay to two photinos. It is easy to see the right hand side of Eq.7 is identically zero.
Figure 1: Diagrams showing (a) the decay of $Z$ to two massless Majorana fermions (left) and (b) the decay of the zino to a photino and a photon (right)
We can extend this even further without much ado to the zero amplitude for a zino going to a photon and a photino, Fig. 1(b). This is inspite of the fact that all three particles are distinct. Only representation theory of the super Poincare algebra along with symmetric/antisymmetric product of irreducible reprsentations provides such a powerful result.
3 Evading the theorem:
In the literature several options have been discussed for evading the theorem through relaxing the conditions for the theorem to hold. For example, one can consider a background of homogeneous electric or magnetic fields [10] there by relaxing the requirement of Lorentz invariance/Poincare symmetry. The external field gives another vector in addition to the polarization vectors and which therefore allows nonzero amplitude for the process. Such background magnetic fields are present in astrophysical contexts. If the massless particles are gluons then there are additional internal color degree of freedom which gives scope for decay of $Z\rightarrow g~{}g$ in QCD[11] (but the caveat is that gluons are always confined).
Much more recently, using noncommutative geometry which is considered as violating Lorentz symmetry, the possibility of $Z\rightarrow\gamma+\gamma$ has been considered. It is well known that massless nature of photon leads to infrared divergences in QED. This was studied extensively in the sixties and was related to the difficulties associated with gauge invariant description of electron which is always surrounded by soft photons (photons of infinitesimal frequencies). This led to a new possibility to evade the Landau-Yang theorem arising from the fact that due to the asymptotic properties of QED the contribution of soft photons can be included in a way which provides Lorentz noninvariant sectors [12].
4 Massive photons and Landau-Yang theorem
Unlike the above cases we consider an alternative which does not require breaking of Lorentz invariance. We will consider the photons to be massive (note that experimentally there is only an upper bound on the mass of the photon, see subsection 4.1 below). It is obvious that if the photons have a small mass the question of the representation theory of the Poincare group takes a new angle. It becomes a question of whether in the tensor product of two massive spin 1 photons obeying Bose symmetry there is a massive spin 1 representation (Z). The answer is obviously yes. This possibility arises because massive photons have three degrees of freedom including the longitudinal photons and Eq.1 is no longer valid. It gives a longitudinal component to the polarization $\vec{s}_{i}$ in the direction of $\vec{k}$ so that $k\cdot\vec{s}_{i}\neq 0$. This allows for a nonzero amplitude for the decay of a massive spin 1 particle like $Z$ to two massive $\gamma$ with a subtantial contribution from the longitudinal component for one of the photons.
4.1 Inonu-Wigner contraction
We mentioned that if the photon has a mass, even if it is small, the Landau-Yang theorem can be evaded. For this one needs to understand the nature of the limit $m_{\gamma}~{}\rightarrow 0$. The question of mass of the photon was raised by Schrodinger himself in 1955 [13] and answered by him that QED results will not be affected by a small photon mass as long as its interaction with matter is only through a conserved current, that is, if $L_{int}~{}=~{}j_{\mu}~{}A^{\mu}$ where $\partial_{\mu}~{}j^{\mu}~{}=~{}0$. Schrodinger himself worked out the upper bound of the mass of the photon using geomagnetic data. The current limit for the mass of the photon as given by the Particle data book [14] is $m_{\gamma}~{}<~{}10^{-18}eV$. The question of mass going to zero limit can be posed as group theoretic question of contraction of $SO(3)~{}\rightarrow E(2)$. The contraction of the massive representation of the Poincare group to the massless representation is discussed in Kim and Wigner [15] and here we briefly summarize the parts relevant for us. The generators of $SO(3)$ can be written as $L_{\pm},L_{3}$ with the Lie-algebra
$\left[L_{3},L_{\pm}\right]~{}=~{}\pm i~{}L_{\pm},\qquad[L_{\pm},L_{\mp}]~{}=~{}2L_{3}$ (8)
On the other hand the group $E(2)$ is specified by $P_{x},P_{y},L_{3}$ with the algebra
$[P_{x},P_{y}]~{}=~{}0,\qquad[L_{3},P_{x,y}]~{}=~{}\pm P_{y,x}$ (9)
The contraction $SO(3)~{}\rightarrow E(2)$ in the limit $m\rightarrow 0$ was provided by Inonu and Wigner through the mapping
$P_{x,y}~{}=~{}\pm S^{-1}L_{x,y}~{}S~{}\qquad S~{}=~{}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&R\end{pmatrix}$ (10)
In the $\epsilon~{}=~{}\frac{1}{R}\longrightarrow 0$ we get $E(2)$. For all values $\epsilon\neq 0$ it provides the same group $SO(3)$. When $\epsilon~{}=~{}0$ we get the sphere of infinte radius close to the northpole and rotations about $x$ and $y$ axes are related to translations along $y$ and $x$ axes, respectively.
This was later escalated to the Poincare group by Kim and Wigner[15]. Both $SO(3)$ and $E(2)$ are subgroups of the Lorentz group and we should arrive at this contraction as part of $SL(2,C)$. This is explained better by looking at the rotation and boost generators $J_{i},~{}K_{i}$ and considering
$N_{1}~{}=~{}K_{1}~{}-~{}J_{2},~{}~{}N_{2}~{}=~{}K_{2}~{}+~{}J_{1}.$ (11)
Then $J_{3},N_{1},N_{2}$ give the Euclidean group $E(2)$. For this we define the light cone coordinates $z_{\pm}~{}=~{}z~{}\pm~{}t$. In the light cone coordinates we define:
$S(R)~{}=~{}e^{-ilog(R)K_{3}}~{}=~{}\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&R&0\\ 0&0&0&\frac{1}{R}\end{pmatrix}$ (12)
where $R~{}=~{}\sqrt{(1+\beta)\over(1-\beta)}$ and $\beta$ is the velocity. The generators $N_{i}$ are proportinal to $N_{i}~{}\propto~{}\frac{1}{R}\epsilon_{ij}~{}S~{}J_{j}~{}S^{-1}$
At the level of the Poincare group this amounts to studying massless representations as the limit of massive representation. Specifically, for the spin 1 particle, the extra degree of freedom that the vector potential gets in the massive case (known as the Stueckelberg degree of freedom) interacts with matter only as a surface term in the massless limit.
5 The process: Stueckelberg Weinberg-Salam model
It is well known that simply adding a mass term to the gauge boson Lagrangian violates the gauge invariance of the theory. To circumvent this we will consider the Stueckelberg extension of the abelian gauge boson in the Weinberg-Salam model [16, 17]. As an explicit example we consider the $Z\rightarrow 2\gamma$ process. $Z$ boson being chargeless is not coupled directly to the photon. We have to look for loop diagrams. Z interacts with the electron and the electron is coupled to the photon. The one loop diagram Fig. 2 is an example of two photon emission from $Z$ (there will be another diagram with $(p,\nu)\leftrightarrow(k,\lambda)$ which we do not show explicitly). In principle we will have contributions from the muon ($\mu$) and tau ($\tau$) loops as well as quark loops. Their contributions are expected to be small.
Figure 2: $Z\rightarrow 2\gamma$ loop diagram
The new amplitude will have an extra term $A_{4}$ in addition to those in Eq.2, where $A_{4}$ is given by:
$A_{4}~{}=~{}\Bigl{(}\vec{s}_{1}\times\vec{s}_{2}\Bigr{)}\cdot\Bigl{(}\vec{S}\times\vec{k}\Bigr{)}$ (13)
This, as mentioned above, is because we no longer have the transversality condition $\vec{k}\cdot\vec{s}_{i}~{}=~{}0$.
Ignoring the coupling and the external gauge bosons, contribution from diagram with only electron in the loop is
$\displaystyle T$$\approx$ $\displaystyle(-1)\int\frac{d^{4}\ell}{(2\pi)^{4}}Tr\left[\gamma^{\mu}\frac{(1-\gamma^{5})}{2}\frac{i}{C_{1}}\gamma^{\lambda}\frac{i}{C_{2}}\gamma^{\nu}\frac{i}{C_{3}}\right]$ (14)
$\displaystyle+(p,\nu)\leftrightarrow(k,\lambda)$
where
[…]
This has to be contracted with the massive gauge boson polarization vectors. As has been pointed out in [17], the Stueckelberg extension of the standard model is anomaly free and therefore we can simply estimate the amplitude for the process without worrying about the details of evaluating the loop integral. Including the coupling constants and the masses of the particles involved, we can estimate $|T|^{2}$ (upto numerical factors) to go as
$|T|^{2}\approx(v^{Z}_{e})^{2}(v^{A}_{e})^{4}M_{Z}^{2}\left(\frac{m_{\gamma}}{m_{e}}\right)^{2}.$ (15)
In the above expression $v^{Z}_{e}$ and $v^{A}_{e}$ are the effective couplings of the electron with the $Z$ boson and the photon (in the notation of [17]), $M_{Z}$ is the mass of the $Z$ boson, $m_{e}$ is the mass of the electron and $m_{\gamma}$ is the mass of the photon  [^3] Here we have projected the $Z-e-e$ vertex to the $V-A$ part since we are only interested in the order of magnitude estimate for the process. However, we note that in the Stueckelberg extension of the standard model there is a slight difference in the electromagnetic coupling of the photon to the left-handed and the right-handed electrons.. We see that in the limit $m_{\gamma}\rightarrow 0$ the $Z\rightarrow\gamma\gamma$ does not occur as required by the Landau-Yang theorem.
Even though for massive photons we have a non-zero contribution, observing such an effect experimentally is difficult because
  1. 1.
    since $m_{\gamma}$ is very small, the probability for the decay $|T|^{2}\propto M_{Z}^{2}(m_{\gamma}/m_{e})^{2}$ is itself very small. Observing such contributions at present is extremely difficult.
  2. 2.
    Observing longitudinal photon by itself is very difficult as its interaction with matter is very small. The only option is the missing energy and momentum carried by it.
In spite of the observational difficulties, a non-zero photon mass, howsoever small, allows for escaping the Landau-Yang theorem without breaking Lorentz/Poincare symmetry which has been the contention of all the other mechanisms. We will comment on issues in connection with the violation of the Landau-Yang theorem in conclusions.
5.1 Stueckelberg extension of SUSY
Lastly, we can extend the above ideas and consider the Stueckelberg extension of supersymmetric QED interacting with a massive vector supermultiplet by providing mass to the photon and its superpartner the photino. For this we need a chiral multiplet in addition to the vector multiplet (for the photon) to cancel the supersymmetric gauge transformations. The chiral multiplet is the supersymmetric analogue of the Stueckelberg field. We refer to the papers for such an extension[9, 18, 19].
6 Conclusions
Landau Yang theorem is an important selection rule which has helped in understanding the role of symmetries in Particle Physics. Since it is related to the massless particles it has acquired importance coming from asymptotic properties of fields. It has also made the understanding of the Hilbert space of states with appropriate properties the central focus. In this connection the mass of the particle as a Lorentz covariant regulator for infrared questions is crucial. Similarly semi-inclusive cross sections for S-matrix in scalar QED, for instance, seem to require mass for the photon as a regulator to remove ambiguities in handling infrared divergences [20]. Furthermore the Landau Yang theorem can also be extended to identical massless fermions. In addition we have also provided extension of the theorem to supersymmetry in this paper.
As Schrodinger[13] pointed out it is important to understand the approach to masslessness of particles even though nature may require zero mass.
Even if we find in Nature the limiting case is realized, we should still feel the urge to adumbrate a theory which agrees with experience on approaching to the limit, not by a sudden. E Schrodinger
A small mass may provide some novel ways of understanding some cosmological questions on the nature of dark matter/energy[21]. Higgs mechanism through which photon can acquire mass may also provide a link between primordial blackholes and dark matter in early galaxies [22, 23]. It is expected that at length scales $L~{}\leq~{}\frac{h}{m_{\gamma}}$ the longitudinal photon behaves like a Stueckelberg scalar. But experimental observation of emission of such a scalar is very difficult if the mass is small as explained by Schrodinger[13]. The only observational signature would be missing energy and momentum in the process $Z\longrightarrow\gamma_{T}~{}+~{}\gamma_{L}$ (which could be a fraction of the energy-momentum carried by the transverse photon, although the rate of such processes will be limited by $M_{Z}(m_{\gamma}/m_{e})^{2}$, as seen earlier).
Acknowledgements: TRG would like to thank Alok Laddha and K. S. Narayan for discussions. Feynman diagrams were made using Feynman Diagram Maker (https://www.aidansean.com/feynman/).